# Anti-Phishing Training Does Not Work: A Large-Scale Empirical Assessment of Multi-Modal Training Grounded in the NIST Phish Scale

### v1

@@ 3.8 Threats to Validity
We encourage the reader to take several important limitations into consideration when interpreting the results of our study. We consider three threats to validity in our work: threats to construct, internal, and external validity [42].
Construct Validity: Construct validity is concerned with whether we operationalized the desired constructs of our study appropriately. In this study, we worked with two primary constructs: (1) the particular interpretations of the two different types of phishing training we deployed, which depend on our cybersecurity training vendor; and (2) our assessment of the difficulty of phishing lures.
It is possible that our cybersecurity vendor’s training was simply lackluster, hence its failure to produce any significant improvement in subject performance. However, note that our results align with prior work that used different trainings. Meanwhile, to control for subjectivity in rating phishing lures, we applied the NIST Phish Scale and used multiple independent analysts to judge difficulty.
Internal Validity: Internal validity threats affect the reliability of conclusions we have about the relationship between the cause and the effect we observe. For our study, we used a stratified random sampling method in order to mitigate the potential biases related to job role and other organizational issues, allowing us to use that as a proxy for other factors such as technical background, educational attainment, etc. Due to the large sample sizes for our two different treatment groups, our study has significant statistical power; however, due to organizational constraints, we were required to limit the sample size for our control group.
External Validity: External threats impact the generalizability of our research. Our study was situated in the USA at a fintech company. The fintech industry context may limit generalizability to other sectors, as financial services employees may have different baseline levels of security awareness when compared to other industries. However, given the minimal effect sizes measured in our study, it is unclear how large of a threat this is. The study’s focus on immediate post-training effects also limits our understanding of long-term training impact. Future research should examine whether the modest effects we observed persist over time or diminish as training recency decreases, or if they improve with volume or other dimensions.
@@ 4 Results
TABLE V: Summary of Training Effectiveness Results. Average rates show that trained groups performed similarly to control group, with modest differences across conditions.
@@
NIST Phish Scale Effect p $<$ 0.001 (highly significant)
@@
Our analysis examined the effectiveness of phishing training across 12,511 participants randomly assigned to three experimental conditions. Table V gives a summary of our key results. Next we present our findings organized by hypothesis, with detailed statistical analysis and effect size reporting for each. Our large-scale study revealed several key findings regarding the effectiveness of phishing training interventions and the predictive utility of the NIST Phish Scale.
@@
4.1 H1: Phish Scale Effect
Training showed mixed effectiveness in our study. While we observed statistically significant effects for some outcomes, effect sizes were generally modest, consistent with recent large-scale findings suggesting limited practical impact of current training approaches [4]. Overall click rate across all participants was 10.4% and overall report rate was 9.6%. Training on both click rates (p = 0.469) and report rates (p = 0.448) were not statistically significant in the primary analysis, though interactions with difficulty levels revealed more nuanced patterns.
The NIST Phish Scale successfully predicted user behavior, supporting H1. Phishing lure difficulty had a highly significant effect on click-through rates (F(2, 12086) = 41.415, p $<$ 0.001, $\eta^{2}$ = 0.007), validating the scale’s utility for measuring phishing complexity. Higher difficulty emails consistently produced higher click rates across all experimental conditions:
@@ 4.1 H1: Phish Scale Effect
This statistically significant difficulty gradient (F(2, 12086) = 41.415, p $<$ 0.001, $\eta^{2}$ = 0.007) confirms that the NIST Phish Scale effectively captures factors that influence user susceptibility to phishing attempts, providing empirical validation for its use in organizational contexts [33, 39].
@@ 4.2 H2: Overall Training Effects
Contrary to H2, training did not produce statistically significant improvements in either click reduction or reporting behavior. Main effect analysis showed no significant difference between trained and untrained groups for click rates (F(2, 12086) = 0.800, p = 0.450) or report rates (F(2, 12086) = 0.874, p = 0.417). This finding aligns with recent large-scale evidence suggesting minimal effectiveness of current training approaches [4].
However, examining treatment effects relative to control group baseline reveals some notable patterns:
Click Rate Changes (relative to control):
  • Training Only: +4.5 percentage points for easy emails, +7.3 for medium, -5.2 for hard
  • Training + Exercise: +3.9 percentage points for easy emails, +5.5 for medium, -5.2 for hard
Report Rate Changes (relative to control):
  • Training Only: +0.1 percentage points for easy emails, +4.7 for medium, +1.9 for hard
  • Training + Exercise: +0.8 percentage points for easy emails, +5.7 for medium, +1.7 for hard
The mixed direction of these effects suggests that training may have differential impacts across difficulty levels, though these did not reach statistical significance in the main effects analysis. The effect of training did not change significantly over time, as we observed similar patterns across the three-month simulation period.
@@ 4.3 H3: Interactive Training Effects
H3 was not supported. Interactive training (Training + Exercise) did not produce significantly higher reporting rates than traditional training alone. While the interactive group showed numerically higher reporting rates across difficulty levels, these differences were not statistically significant. The coefficient estimate for Training + Exercise relative to Training Only was modest and non-significant across all difficulty levels.
@@ 4.4 H4: Training-Difficulty Interactions
H4 was not supported. We found no statistically significant interaction between training modality and lure difficulty for click rates (F(4, 12086) = 3.135, p = 0.014) or report rates (F(4, 12086) = 0.517, p = 0.723). While there was a marginal interaction effect for click rates, this primarily reflected the pattern that training appeared to increase clicks on easier emails while potentially reducing clicks on harder emails.
@@
  • Training Only × Hard: -9.67 percentage points (p = 0.002)
  • Training + Exercise × Hard: -9.09 percentage points (p = 0.004)
@@ 4.5 Effect Sizes and Practical Significance
While some individual coefficients reached statistical significance, effect sizes were generally small ($\eta^{2}$ $<$ 0.01 for all main effects), raising questions about practical significance [7]. The largest meaningful effect was the difficulty gradient ($\eta^{2}$ = 0.007 for click rates), confirming the NIST Phish Scale’s predictive validity.
@@
Figure 5: Comparison of effect sizes ($\eta^{2}$) for key factors in phishing susceptibility. The NIST Phish Scale (difficulty) shows the strongest and only practically meaningful effect ($\eta^{2}=0.007$, $p<0.001$), while training and interaction effects are negligible ($\eta^{2}<0.01$).
4.6 Other Measurements
The final question we considered was whether there was any relationship between training time and detection ability, e.g., decay of training effectiveness. In our study, subjects received the phishing simulation email between 0–4 months after the training. Figure 6 shows the click rate by treatment group, discriminating by lure difficulty. We observe no discernible trend in the results as a function of time. Given the result of the visual inspection, we did not perform statistical tests of this aspect.
@@ 5 Discussion
Our large-scale field study provides empirical evidence for both the utility of the NIST Phish Scale and the limited effectiveness of current phishing training approaches. While we observed some statistically significant effects, the practical significance of these findings raises important questions about the return on investment for organizational security awareness programs.
5.1 Recommendations for Practice
@@
Confronting Training Effectiveness Realities: Despite deploying well-structured training programs to over 12,000 employees, we found no statistically significant main effects of training on either click rates (p = 0.450) or report rates (p = 0.417). The small effect sizes we observed ($\eta^{2}$ $<$ .01 for all main effects) indicate that even statistically significant improvements translate to minimal operational impact. This aligns with recent large-scale evidence from Ho et al. [4] and challenges the optimistic assessments often promoted by cybersecurity training vendors.
The disconnect between vendor marketing claims and empirical evidence suggests that the substantial organizational resources typically invested in security awareness programs may not justify the modest improvements observed—even in best-case scenarios. Organizations must acknowledge these limitations to make informed investment decisions about cybersecurity resources.
@@
Layered Defense: Our findings necessitate a fundamental shift in organizational phishing defense strategy. Rather than viewing training as a primary protection mechanism, organizations should treat security awareness programs as one component of a comprehensive layered defense strategy. The strong effect of phishing difficulty on user susceptibility (F(2, 12086) = 41.415, p $<$ 0.001) demonstrates that sophisticated attacks will likely succeed regardless of training efforts.
This evidence suggests several strategic reorientations: prioritizing passwordless authentication systems using passkeys and FIDO-compliant hardware tokens like YubiKeys, especially for administrative roles; implementing comprehensive password manager deployment to reduce credential reuse vulnerabilities; enhancing email security controls specifically designed to catch sophisticated attacks; and developing incident response procedures that assume successful phishing will occur despite training efforts.
The reporting behavior improvements we observed, while modest, indicate that interactive training components may have some value in encouraging security-conscious behavior. However, organizations must weigh the practical significance of 1-2 percentage point improvements against the substantial additional resources required for interactive training delivery.
Implementation Framework: Organizations seeking to optimize their phishing defense should implement the following evidence-based framework:
Technical Controls Priority. Invest primarily in email filtering, attachment analysis, URL protection, and multi-factor authentication systems. Our findings suggest these provide more reliable protection than user training against sophisticated attacks.
Targeted Training. Rather than broad-based programs, implement focused training for specific roles, departments, or threat scenarios where evidence suggests training may be more effective. Use standardized difficulty assessment frameworks like the NIST Phish Scale for meaningful evaluation.
@@
Continuous Assessment. Regularly evaluate the cost-effectiveness of training programs against alternative security investments, using standardized metrics that capture both behavioral outcomes and resource allocation efficiency.
This approach recognizes the sobering reality that current training approaches provide only modest protection against evolving phishing threats. The goal is not to eliminate training but to right-size its role within a comprehensive security strategy informed by empirical evidence rather than vendor marketing claims.
5.2 NIST Phish Scale Validation and Practical Utility
Our study provides strong empirical validation for the NIST Phish Scale’s predictive utility in organizational contexts. The highly significant effect of lure difficulty on click-through rates (F(2, 12086) = 41.415, p $<$ 0.001) demonstrates that the scale effectively captures factors influencing user susceptibility to phishing attempts. The clear gradient in click rates across difficulty levels—from 7.0% for easy emails to 15.0% for hard emails—confirms that the scale’s two-dimensional framework meaningfully differentiates phishing sophistication.
This validation extends the early work of Barrientos et al. [39] by demonstrating the scale’s effectiveness in a large operational environment. Unlike previous smaller-scale validations, our study shows that the NIST Phish Scale maintains its predictive power across diverse organizational roles, departments, and user populations. This supports the scale’s adoption as a standardized framework for evaluating phishing simulation programs and benchmarking organizational vulnerability assessments.
However, the scale’s practical implementation revealed important considerations for practitioners. The process of systematically rating premise alignment across thousands of employees proved more complex than anticipated, requiring careful consideration of organizational communication patterns and role-specific email environments. Future implementations should account for these practical challenges when scaling the framework to enterprise environments.
5.3 Training Modality Effects and Interaction Patterns
While interactive training did not significantly outperform traditional lecture-based training in our primary analysis, the interaction patterns we observed provide nuanced insights into training effectiveness. The significant negative interactions between both training conditions and hard difficulty emails (Training Only × Hard: -9.67 percentage points, p = 0.002; Training + Exercise × Hard: -9.09 percentage points, p = 0.004) suggest that training may provide some protection against the most sophisticated phishing attempts. Conversely, the positive coefficients for training conditions with easier emails indicate that training may inadvertently increase vulnerability to simpler phishing attempts. This counter-intuitive finding aligns with concerns raised in prior work about potential negative effects of training programs [23]. One possible explanation is that training focuses employees’ attention on detecting sophisticated attacks while reducing their vigilance toward obviously fraudulent emails.
These patterns underscore the complexity of human factors in cybersecurity training. Rather than uniformly improving detection across all threat types, training appears to redistribute user attention and response patterns in ways that may create new vulnerabilities. This highlights the importance of comprehensive evaluation frameworks that examine training effects across varied attack sophistication levels.
5.4 The Evolving Threat Landscape and Future Challenges
@@
Our findings must be interpreted within the context of rapidly evolving phishing threats. The emergence of large language models and generative AI technologies is enabling attackers to create increasingly sophisticated and personalized phishing attempts that may render current training approaches obsolete. The traditional approach of teaching users to identify visual cues and suspicious content becomes less effective when attackers can generate near-perfect replications of legitimate communications.
This technological evolution suggests that the cybersecurity community may need to fundamentally reconsider human-centered defense strategies. Rather than expecting users to reliably detect sophisticated attacks, future approaches might focus on cryptographic verification systems, zero-trust architectures, in-person interaction (supporting return-to-office/RTO mandates) or other technical mechanisms that reduce reliance on human judgment.
The broader threat model is also expanding beyond traditional email phishing to include SMS phishing (smishing), voice phishing (vishing), and social engineering attacks using deep fakes across multiple communication channels. Training programs that focus narrowly on email-based threats may provide limited protection against this diversified attack landscape.
5.5 Implications for Theory and Practice
@@
Our findings have significant implications for both cybersecurity theory and organizational practice. The limited training effectiveness we observed suggests that current human-centered defense paradigms may be fundamentally inadequate for addressing sophisticated phishing threats. The modest effect sizes ($\eta^{2}$ $<$ 0.01 for all main effects) indicate that even statistically significant improvements may not translate to meaningful operational risk reduction.
@@
The interaction patterns we identified where training appeared to increase vulnerability to easier emails while potentially providing some protection against harder emails highlight the complex and sometimes counterintuitive effects of security awareness interventions. These findings suggest that training programs may redistribute rather than uniformly reduce user vulnerability, creating new attack vectors while addressing others.
@@
5.6 The Future of Human-Centered Cybersecurity
Our findings emerge at a particularly critical juncture in the evolution of cybersecurity. With the rapid advancement of generative AI enabled by large language models and other technologies, attackers are enabled to create increasingly sophisticated and personalized phishing attempts that may render traditional cue-based detection obsolete. When attackers can automate the creation of personalized phishing lures and generate near-perfect replicas of legitimate communications, it becomes unreasonable to expect users to reliably identify deceptive content.
The fundamental challenge lies in the inherent complexity of human factors in resistance to phishing attacks. Since these exploits target individual vulnerabilities, personal and contextual factors play an enormous role in individual user susceptibility [19]. Even a single individual will vary dramatically throughout a day—users typically focus on content without paying attention to security, and are significantly impacted by cognitive constraints such as memory limitations and current cognitive load, all of which affect vigilance [43, 44]. Research has investigated variability in individual attentiveness based on factors like gender, age, and technical expertise, though the literature remains mixed [45, 23, 38]. Training represents only a single layer of defense against these multifaceted human vulnerabilities.
This technological evolution, combined with the complexity of human decision-making under cognitive load, suggests that the cybersecurity community must fundamentally reconsider where human-centered defenses lie in the multilayered approach organizations must take to cybersecurity. Future approaches may need to emphasize cryptographic verification systems, zero-trust architectures, and other technical mechanisms that can reduce reliance on the human firewall rather than attempting to enhance and upgrade the human firewall, which runs on hardware with a broad spectrum of variability.
The expanding threat landscape—encompassing SMS phishing, voice phishing, and multi-channel social engineering attacks—further challenges the viability of training-centric approaches focused primarily on email-based threats. As attack vectors diversify and sophistication increases, the cognitive burden placed on users to maintain vigilance across all communication channels may exceed reasonable human capabilities, particularly given the natural fluctuations in attention and decision-making quality that characterize normal human cognition. The role of human factors in resistance to phishing is complex.

### v2

@@ 3.6 Threats to Validity
We encourage the reader to consider several limitations when interpreting the results of our study. We distinguish these limitations into three kinds of threats to validity. Following the guidance of Verdecchia et al., we focus on what we perceive to be the most crucial ones [94].
Construct Validity: Construct validity is concerned with whether we operationalized the desired constructs of our study appropriately [84, 98]. In this study, we worked with two primary constructs: (1) the particular interpretations of the two different types of phishing training we deployed, which depend on our cybersecurity training vendor; and (2) our assessment of the difficulty of phishing lures.
It is possible that our cybersecurity vendor’s training was simply lackluster, hence its failure to produce any significant improvement in subject performance. However, note that our results align with prior work that used different trainings. Meanwhile, to control for subjectivity in rating phishing lures, we applied the NIST Phish Scale and used multiple independent analysts to judge difficulty [85].
Internal Validity: Internal validity threats affect the reliability of conclusions we have about the relationship between the cause and the effect we observe [84, 98]. For our study, we used stratified random sampling to mitigate potential biases related to job role and other organizational issues. Due to the large sample sizes for Treatments A and B, our study has significant statistical power. Organizational constraints required us to limit the sample size for our control group. This group was still large enough to provide a baseline.
Like all real-world phishing measurement studies, our work faces several measurement concerns. False positive clicks may occur when automated security systems (security gateways, antivirus, etc.), shadow IT applications, or personally installed tools / browser extensions automatically scan or prefetch URLs from emails, generating click events without actual user interaction. False negative reporting may also occur, wherein users report phishing emails through alternative channels, such as contacting the IT helpdesk directly, rather than the simulation platform’s designated “Report Phish” button. These behaviors would, respectively, incorrectly increase and decrease our measurements of subject interactions. We cannot fully eliminate such threats, but our large sample size mitigates their impact on our overall conclusions.
External Validity: External threats impact the generalizability of our research [81, 100]. Our study was situated in the USA at a fintech company. The fintech industry context may limit generalizability to other sectors, as financial services employees may have different baseline levels of security awareness when compared to other industries. However, given the minimal effect sizes measured in our study, it is unclear how large of a threat this is. The study’s focus on immediate post-training effects also limits our understanding of long-term training impact. Future research should examine whether the modest effects we observed persist over time or diminish as training recency decreases, or if they improve with volume or other dimensions.
@@ 4 Results
Table 5: Summary of Training Effectiveness Results. Average rates show that trained groups performed similarly to control group, with modest differences across conditions. Effect sizes are shown in Figure˜5.
@@
NIST Phish Scale Effect p << 0.001 (highly significant)
@@
Our analysis examined the effectiveness of phishing training across 12,511 participants randomly assigned to three experimental conditions. Table˜5 gives a summary of our key results. Figure˜5 complements this with a depiction of effect sizes for the controlled factors.
Figure 5: Comparison of effect sizes ($\eta^{2}$) for key factors in phishing susceptibility. The NIST Phish Scale (difficulty) shows the strongest and only practically meaningful effect ($\eta^{2}=0.007$, $p<0.001$), while training and interaction effects are negligible ($\eta^{2}\leq 0.001$). See also Table˜5.
4.1 H1: Phish Scale Effect
@@
H1 was supported. Figure˜6 illustrates the comparison of effect sizes across different factors in phishing susceptibility. The NIST Phish Scale successfully predicted user behavior. Phishing lure difficulty had a highly significant effect on click-through rates (F(2, 12086) = 41.415, p << 0.001, $\eta^{2}$ = 0.007), validating the scale’s utility for measuring phishing complexity. Higher difficulty emails consistently produced higher click rates across all experimental conditions:
@@
This statistically significant difficulty gradient confirms that the NIST Phish Scale effectively captures factors that influence user susceptibility to phishing attempts, providing empirical validation for its use in organizational contexts.
@@ 4.2 H2: Overall Training Effects
H2 was not supported. Training did not produce statistically significant improvements in either click reduction or reporting behavior. Main effect analysis showed no significant difference between trained and untrained groups for click rates (F(2, 12086) = 0.800, pp=0.450) or report rates (F(2, 12086) = 0.874, pp=0.417). This finding aligns with recent large-scale evidence suggesting minimal effectiveness of current training approaches [48].
@@ 4.3 H3: Interactive Training Effects
H3 was not supported. Treatment B, interactive training (Training + Exercise) did not produce significantly higher reporting rates than traditional training alone. While the interactive group showed numerically higher reporting rates across difficulty levels, these differences were not statistically significant. The coefficient estimate for Training + Exercise relative to Training Only was modest and non-significant across all difficulty levels.
@@ 4.4 H4: Training-Difficulty Interactions
H4 was not supported. We found no statistically significant interaction between training modality and lure difficulty for click rates (F(4, 12086) = 3.135, pp=0.014) or report rates (F(4, 12086) = 0.517, pp=0.723). While there was a marginal interaction effect for click rates, this primarily reflected the pattern that training appeared to increase clicks on easier emails while potentially reducing clicks on harder emails.
@@
  • Training Only × Hard: -9.67 percentage points (pp=0.002)
  • Training + Exercise × Hard: -9.09 percentage points (pp=0.004)
@@
4.5 H5: Reporting Timeliness
H5 was partially supported, though with important caveats about sample size limitations. Our analysis of reporting timeliness revealed mixed patterns in organizational-level temporal protection across treatment groups. Individual reporting behavior showed a median time-to-report of 21 minutes (0.35 hours), with 90% of reports occurring within 18.8 hours of phishing deployment, indicating generally rapid user response to suspicious emails.
At the organizational level, we observed notable differences in inoculation patterns—campaigns where reports preceded clicks, enabling organizational threat mitigation before exploitation. Training Only achieved the highest inoculation rate at 54.5% of campaigns, with a median organizational response time of -0.33 hours (reports preceding clicks). In contrast, the Control group achieved inoculation in 42.9% of campaigns with a median response time of 8.09 hours, while Training + Exercise showed 36.4% inoculation with a median response time of 1.21 hours.
However, these differences were not statistically significant due to necessarily small sample sizes at the campaign level (n=33 campaigns total). The Organizational Inoculation Index revealed substantial organizational resilience independent of individual training effectiveness, with over one-third of all template-group combinations demonstrating proactive reporting patterns. While Training Only showed numerically superior temporal performance, the limited campaign-level sample size prevents definitive conclusions about training effects on organizational-level reporting timeliness.
These findings provide the first empirical measurement of organizational temporal security dynamics, directly addressing Steves et al.’s research question about whether "time to first report is sooner than time to first click" [88]. The results suggest that organizational-level protective behaviors may operate independently of individual training effectiveness, though larger-scale studies are needed to establish statistical significance of these temporal patterns.
4.6 Other Measurements
The final question we considered was whether there was any relationship between training time and detection ability, e.g., decay of training effectiveness. In our study, subjects received the phishing simulation email between 0–4 months after the training. Figure˜7 shows the click rate by treatment group, discriminating by lure difficulty. We observe no discernible trend in the results as a function of time. Given the result of the visual inspection, we did not perform statistical tests of this aspect.
@@ 4.7 Effect Sizes and Practical Significance
While some individual coefficients reached statistical significance, effect sizes were generally small ($\eta^{2}$ << 0.01 for all main effects), raising questions about practical significance [90]. The largest meaningful effect was the difficulty gradient ($\eta^{2}$ = 0.007 for click rates), confirming the NIST Phish Scale’s predictive validity.
@@ 5 Discussion
Our large-scale field study provides empirical evidence for both the utility of the NIST Phish Scale and the limited effectiveness of current phishing training approaches. While we observed some statistically significant effects, their practical significance raises important questions about the return on investment for organizational security awareness programs.
5.1 The Reality of Training (In)Effectiveness and Organizational Security
@@
Confronting Training Effectiveness Realities: Despite deploying well-structured training programs to over 12,000 employees, we found no statistically significant main effects of training on either click rates (pp=0.450) or report rates (pp=0.417). The small effect sizes we observed ($\eta^{2}$ < 0.01 for all main effects) indicate that even statistically significant improvements translate to minimal operational impact. This aligns with recent large-scale evidence from Ho et al. [48] and challenges the optimistic assessments often promoted by cybersecurity training vendors.
Our findings suggest that many organizations may be operating under a dangerous illusion of training effectiveness. The cybersecurity community must confront this reality by demanding more rigorous evaluation standards and refusing to accept vendor claims that cannot be substantiated under realistic threat conditions. Only by acknowledging the true limitations of current training approaches can organizations make informed decisions about resource allocation and security strategy [37].
The Possibility of Training Metrics Manipulation: Our findings suggest the possibility that organizations may inadvertently—or deliberately—manipulate phishing training metrics to achieve desired outcomes. The strong predictive power of the NIST Phish Scale (F(2, 12086) = 41.415, p < 0.001) demonstrates that practitioners can significantly influence training results simply by adjusting the difficulty of phishing lures used in simulations.
Our observation informs broader concerns about “compliance theater” in cybersecurity [44], where organizations optimize for metrics that satisfy regulatory requirements rather than actual security improvement. Recent work by Ho et al. [48] and Lain et al. [53, 54] provides additional evidence that training effectiveness is often overstated. Our findings demonstrate a specific mechanism through which such overstatement may occur: the systematic use of phishing lures that fail to represent realistic attack sophistication.
Such selection may not be the fault of the organizations alone. We note that many vendors of phishing training allow organizations to select from libraries of phishing templates without standardized difficulty assessment. This enables the selection—accidental or deliberate—of easier lures, resulting in inflated success metrics [18].
5.2 Recommendations for Phish Resilience
@@
Layered Defense Strategy: Our findings add to the evidence that organizational phishing defense strategy must treat training as one of many layers of defense. The strong effect of phishing difficulty on user susceptibility (F(2,12086) = 41.415, p < 0.001) indicates that attacks are likely to succeed regardless of training efforts.
Evidence-Based Implementation Framework: Organizations seeking meaningful security assessment must resist the temptation to achieve favorable metrics through lure selection bias. Our results support the following three-part framework:
(1) Technical controls first. As humans cannot be trained to reliably detect phishing, automated defenses are in order. Prior work recommends investing primarily in filtering, attachment analysis, URL protection, and multi-factor or phishing-resistant authentication [4].
(2) Targeted training. Our results suggest that generic company-wide training is ineffective [48]. Prefer focused, role- or scenario-specific training where prior evidence suggests impact [82, 90]. Use standardized difficulty frameworks (e.g., the NIST Phish Scale) for evaluation.
@@
(3) Continuous assessment. Regularly compare training costs and outcomes against alternative investments using standardized metrics that reflect behavior and resource efficiency.
In phishing training assessment, guided by our methods, we recommend that organizations:
  • Use standardized difficulty frameworks like the NIST Phish Scale to ensure representative threat simulation
  • Report results separately by difficulty level rather than aggregating across lure types
  • Benchmark against external standards rather than setting arbitrary internal targets
  • Focus on improvement trends over time rather than absolute performance thresholds
We do not recommend eliminating phishing training, but rather to right-size its role within a defense-in-depth strategy guided by evidence rather than marketing claims.
@@
The Evolving Threat Landscape: Our findings must be interpreted within the context of rapidly evolving phishing threats. The emergence of large language models and generative AI technologies is enabling attackers to create increasingly sophisticated and personalized phishing attempts that may render current training approaches obsolete [4, 80, 9]. When attackers can automate the creation of personalized phishing lures and generate near-perfect replicas of legitimate communications, it may become unreasonable to expect users to reliably identify deceptive content. [^5] One of the authors recently failed his organization’s annual phishing training, which was experimenting with GenAI-based lures.
The fundamental challenge lies in the inherent complexity of human factors in resistance to phishing attacks. Since these exploits target individual vulnerabilities, personal and contextual factors play an enormous role in individual user susceptibility [30]. Even a single individual will vary dramatically throughout a day—users typically focus on content without paying attention to security, and are significantly impacted by cognitive constraints such as memory limitations and current cognitive load [33, 32]. Training represents only a single layer of defense against these multifaceted human vulnerabilities.
This technological evolution, combined with the complexity of human decision-making under cognitive load, suggests that the cybersecurity operations community must reconsider where human-centered defenses lie in the multilayered approach organizations must take to cybersecurity. Future approaches may need to emphasize cryptographic verification systems, zero-trust architectures, and other technical mechanisms to reduce reliance on the “human firewall”.
@@
Implications for Theory and Practice: Our findings have significant implications for both cybersecurity theory and organizational practice. The limited training effectiveness we observed suggests that current human-centered defense paradigms may be fundamentally inadequate for addressing sophisticated phishing threats. The modest effect sizes ($\eta^{2}$ < 0.01 for all main effects) indicate that even statistically significant improvements may not translate to meaningful operational risk reduction.
@@
The interaction patterns we identified—where training appeared to increase vulnerability to easier emails while potentially providing some protection against harder emails—highlight the complex and sometimes counterintuitive effects of security awareness interventions. These findings suggest that training programs may redistribute rather than uniformly reduce user vulnerability, creating new attack vectors while addressing others.
@@
5.3 Future Work
Our findings open several important avenues for future research that could advance understanding of phishing training effectiveness and organizational cybersecurity resilience.
AI-Generated Phishing and the Limits of the NIST Phish Scale: The emergence of generative AI enables highly convincing phishing attacks that may lack the flaws traditionally used to assess difficulty, such as poor grammar or implausible scenarios [80]. This calls into question the continued relevance of the NIST Phish Scale, which assumes such imperfections are present. Although some phishing campaigns will continue to use low-budget approaches with human-induced errors, future research should investigate whether training methods can prepare users to recognize AI-crafted phishing attempts and whether new paradigms are needed to simulate and defend against these advanced threats.
Automated Difficulty Assessment: The manual process of rating phishing lures using the NIST Phish Scale is subject to rater subjectivity and does not scale well to large deployments. Future work should explore automated or semi-automated systems—potentially powered by machine learning—to assess lure difficulty by evaluating features such as cue visibility and premise plausibility. Additionally, comparing lure success rates in randomized control groups with sufficiently large populations could offer a data-driven alternative for estimating difficulty. These approaches could improve consistency, scalability, and real-world applicability of phishing difficulty assessments.
Alternative Training Modalities and Interventions: While our study focused on lecture-based and interactive training approaches commonly offered by vendors, other intervention types remain understudied. Future research should investigate the effectiveness of just-in-time training, personalized adaptive tools that adjust to users’ individual risk profiles [82], and hybrid models that combine human training with technical controls. Emerging interventions—such as AI-driven, real-time reviews of links and emails—also warrant exploration. Even widely deployed tools like spelling and grammar checkers may offer untapped potential as passive defenses against phishing.
Organizational Inoculation Index Refinement: Our introduction of the Organizational Inoculation Index offers a starting point for measuring collective security behaviors across a workforce, but several refinements could improve its practical value. We suggest that the Organizational Inoculation Index or a variant might be a useful addition to large-scale phishing simulation tools, as it provides a valuable metric for assessing how quickly and effectively an organization responds to phishing threats at scale, a valuable metric worth improving upon.
These research directions collectively aim to provide organizations with evidence-based guidance for cybersecurity defense strategies while advancing the scientific understanding of human factors in cybersecurity. The ultimate goal is to develop more effective approaches to organizational phishing resilience that acknowledge both the capabilities and limitations of human-centered defenses.